Write molecular equation, ionic equation, net ionic equation for the following reactions

1. reaction of AgNO3 and NaCl

2. reaction of AgNO3 and NaOH

3. reaction of AgNO3 and Na2SO4

4. reaction of AgNO3 and Na3PO4

5. reaction of AgNO3 and Na2CO3

6. reaction of AgNO3 and NaC2H3O2
Answer #1

1.AgNO3 + NaCl = AgCl + NaNO3

2.AgNO3 + NaOH = AgOH + NaNO3

3.AgNO3 + Na2SO4 = Ag2SO4 + NaNO3

4.AgNO3 + Na3PO4 = Ag3PO4 + NaNO3

5.AgNO3 + Na2CO3 = Ag2CO3 + NaNO3

6.AgNO3 + NaC2H3O2 = AgC2H3O2 + NaNO3